# Mechanical Properties of the Proton from a Deformed AdS Holographic Model

### v1

@@ III.1 $A(q^{2})$ form factor
Using the numerical solutions of (14) and (22), we numerically solve the integral in (24) as a function of $q^{2}$. In order to compare it with the lattice QCD results, we consider a normalization of our numerical $A(q^{2})$ by $A(0)$ of the lattice data, since this initial value is not fixed in holography. Using an analytical soft-wall model, the authors in Mamo:2019mka proceeded by choosing $A(0)$ from the lattice data of Shanahan:2018pib to fix their corresponding zero momentum squared value and adjust their analytical result to these data. This can be achieved by considering a target value, say $A(0)_{\text{targ}}$ from the lattice, and then a normalized curve given by
@@
Here we choose to compare our holographic model with the more up to date Lattice data of Hackett:2023rif related to the gluon contribution to $A(q^{2})$. In this work, however, we extend the normalization procedure described above by choosing a range of target $A(0)_{\text{targ}}$ values that comprises the zero momentum values of Shanahan:2018pib and Hackett:2023rif, as well as their error bars. We choose this range as $A(0)_{\text{targ}}=0.4-0.7$. Then, we perform a $\chi^{2}$ minimization procedure as follows. The following $\chi^{2}$ functional form is given by
$\chi^{2}=\sum_{i=2}^{N}\frac{[A(q^{2}_{i})_{\text{norm}}-A(q^{2}_{i})_{\text{lat}})]^{2}}{\sigma^{2}}$ (26)
where $i=2$ corresponds to the second value of both the normalized holographic model and the lattice value, since they are built to be equal and hence give $\chi^{2}=0$. $N$ represents the total number of momenta transferred. $\sigma=0.5\,(A_{\text{upper}}-A_{\text{lower}})$ corresponds to the symmetric error associated with the upper and lower points of the error bars of the data from Hackett:2023rif. Keeping all other parameters in the model fixed, we look for the value of $A(0)_{\text{targ}}$ in $A(q^{2}_{i})_{\text{norm}}$ that minimizes (26). In this context, we find $A(0)_{\text{targ}}=0.628$ to be the best target value found. We point out that this target value was entered as the only degree of freedom used in this minimization procedure. This gives $\chi^{2}_{\text{min}}=32.472$ and $\chi^{2}/\text{ndf}=0.984$. These results give the best-fit curve of our normalized holographic model, as illustrated in Fig. 1.
Figure 1: $A(q^{2})_{\text{norm}}$ form factor adjusted to the gluon contributions of lattice data of Shanahan:2018pib and Hackett:2023rif. The green dots correspond to the gluonic contribution in lattice QCD of Hackett:2023rif, and the blue curve represents our best-fit normalized holographic form factor AA .The purple curve represents the soft-wall result of Mamo & Zahed Mamo:2019mka, compared to the lattice data of Shanahan:2018pib (red dots).
The gluonic radius related to the momentum distribution is given by Polyakov:2018zvc
@@ III.2 $D(q^{2})$ form factor
From the numerical integration of (28), one can calculate the form factor $D$ defined by $D(q^{2})\equiv 4\,C(q^{2})$. We perform the same normalization and $\chi^{2}$ minimization procedure for $D(q^{2})$. Similarly to $A(0)_{\text{targ}}$, we choose to restrict the target $D(0)$ in a normalized $D(q^{2})_{\text{norm}}$ to the range $D(0)_{\text{targ}}=(-1.5)-0.0$, in which most of the lattice data from Hackett:2023rif lie. The resulting best target value $D(0)_{\text{targ}}$ is found to be $D(0)_{\text{targ}}=-1.192$. The resulting fit to the lattice data of Hackett:2023rif is illustrated in Fig. 2. We can notice that we get close to the QCD data of the lattice, but the fit is not as good as that achieved for the form factor $A$, especially for the high momenta transferred $q^{2}$. Indeed, the minimization procedure described above, when applied here to $D(q^{2})_{\text{norm}}$ with a target $D(0)_{\text{targ}}$, gives $\chi^{2}_{\text{min}}=64.219$ with $\chi^{2}/\text{ndf}=4.281$, which quantitatively states that the model deviates a considerable amount from the data considered. In fact, this seems to be a common feature shared between bottom-up holographic models such as the one used in this work and the soft-wall model considered in Mamo:2019mka. However, even though there is this discrepancy between our holographic model and the data considered for comparison, we can still explore the fact that its trend qualitatively follows the gluon lattice data of Shanahan:2018pib and Hackett:2023rif to study the mechanical properties of the proton, such as the pressure and shear distributions in this model.
Figure 2: Holographic $D(q^{2})_{\text{norm}}$ form factor from deformed AdS model (blue curve) compared to gluon lattice data of Hackett:2023rif (green dots). The purple curve corresponds to the soft-wall model of Mamo:2019mka, which compares to the lattice data of Shanahan:2018pib (red dots).
@@ IV Pressure, Shear, and Energy Distributions
This is because these distributions can be formulated in terms of $\tilde{D}(r)$ according to Polyakov:2018zvc
@@
with $m=M_{p}$ the proton mass. For numerical convenience, instead of using the numerical result for $D(q^{2})_{\text{norm}}$, we use a dipole approximation to it, which is well behaved when computing its Fourier transform. We consider
$D(q^{2})=\frac{D_{0}}{\left(1+\frac{q^{2}}{\Lambda^{2}}\right)^{2}},$ (32)
where $D_{0}$ and $\Lambda$ are free parameters. A very good fit to our numerical result $D(q^{2})_{\text{norm}}$ is found with $D_{0}=-1.196$ and $\Lambda=1.165\,\text{GeV}$. The pressure and shear distributions are shown by Figures 4 and 4.
Figure 3: Pressure distribution from the dipole approximation (32), with a repulsive (red) and confining (green) pressure regions.
Figure 4: Shear distribution from the dipole approximation (32).
One can also obtain the mechanical radius associated to those distributions, given by Polyakov:2018zvc
$\langle\,r_{\text{mech}}^{2}\rangle=\frac{6}{\Lambda^{2}}=(0.42\,\text{fm})^{2}.$ (33)
These results are comparable with experimental results in Burkert:2018bqq and lattice QCD ones, as in Shanahan:2018nnv.
One can notice from these predictions that the strong force behaves as a repulsive force in the inner regions of the proton, whereas it behaves as a confining force in its outer regions. These positive and negative pressure regions present in Fig. 4 integrate to $2.31\times 10^{-6}\,\text{GeV}$, which, within the precision of our calculation, is numerically consistent with the von Laue stability condition laue1911dynamik
@@
$\epsilon(r)=m\,\left[A(q^{2})+\frac{q^{2}(D(q^{2})+A(q^{2})-2\,B(q^{2}))}{4\,m^{2}}\right]_{\text{FT}}.$ (35)
where FT stands for Fourier transform. Here, the $A$ and $D$ form factors represent our normalized ones obtained in III and, in our case, $B=0$. We also need to obtain an approximation to our numerical $A(q^{2})_{\text{norm}}$ to compute the Fourier transform above, and again consider a dipole parameterization of the form
@@
with $A_{0}$ and $\lambda$ free parameters. The best approximation to our numerical $A(q^{2})_{\text{norm}}$ is found with $A_{0}=0.627$ and $\lambda=1.186\,\text{GeV}$. Hence, the resulting energy density can be computed and is displayed in Fig. 5
Figure 5: Energy distribution from the dipole approximations (32) and (36).
This energy distribution is very similar to that obtained in lattice QCD by Hackett:2023rif.
@@ V Normal and Tangential Forces
Following Polyakov:2018zvc, we consider a slice of the proton defined in spherical coordinates by $\theta=\pi/2$. The forces acting on the infinitesimal area element $d\textbf{S}=dS_{n}\,\textbf{e}_{n}+dS_{\theta}\,\textbf{e}_{\theta}+dS_{\phi}\,\textbf{e}_{\phi}+$ of the slice are
@@
Figure 6: Normal and tangential forces on a spherical shell of radius $r$ in the nucleon.
@@
Figure 7: Force field visualization of the tangential forces $F_{\theta}$ and $F_{\phi}$ as functions of $r$. The left panel displays the full profile of these forces. The right panel zooms in into the inner region where the sign of the forces changes.
Figure 8: Force field visualization of the normal force $F_{n}(r)$ inside the proton.
@@ VI Conclusions
In the present work, we employ the deformed AdS background holographic model to describe gravitational form factors of the proton and explore some of its mechanical properties based on the $D$ term, focusing on the contribution of the gluon to them. We find a very good agreement between our prediction for the gluonic form factor $A$ and the recent lattice QCD computation performed in Hackett:2023rif, using a normalization procedure to fit $A(0)$ to the lattice result with a target $A(0)_{\text{targ}}$ that minimizes the $\chi^{2}$ measure of the difference between our model prediction and the lattice data. Its trend is also compared to the previous holographic result of Mamo:2019mka as well as to the previous lattice data of Shanahan:2018pib. The same approach is carried out for the gluonic term $D$, resulting, in this case, in a qualitative agreement between our model and the lattice data of Hackett:2023rif, whose behavior draws a particular shape not fully captured by the holographic model used in this work. However, we remark that those results were obtained numerically in all the steps present in their calculations, in contrast to previous holographic results using the soft-wall model in Abidin:2009hr; Mamo:2019mka, which use analytical solutions to form factors. The general agreement between the numerical results in our work, which are also in the ballpark of the results present in Shanahan:2018pib; Mamo:2019mka, and the lattice results of Hackett:2023rif, shows that this model is capable of capturing features present in both the lattice QCD literature and previous holographic models.
@@
From the $D$ term, since its trend follows the behavior present in the lattice results, we can explore how the model describes some mechanical properties of the nucleon. Using a dipole approximation to our numerical result, which is better suited for the calculation of the Fourier transform of $D(r)$ and $A(r)$, we can compute the pressure and shear distributions inside the proton, which balance each other to keep the system stable. This is supported by the small value of the integral $\int r^{2}\,p(r)\,dr=2.31\times 10^{-6}\,\text{GeV}$ in our approximation, which should be zero in the exact case. We are also able to evaluate the energy density in the proton using a dipole approximation to $A(q^{2})_{\text{norm}}$. The results follow similar trends present in the literature, e.g., such as those in Lorce:2018egm; Polyakov:2018zvc; Shanahan:2018nnv; Burkert:2018bqq; Burkert:2021ith; Mamo:2019mka; Mamo:2021krl. However, the mechanical radius associated with these distributions is smaller than the ones found in these works. From pressure and shear distributions, internal forces are also investigated, presenting a similar behavior appearing in this literature.

### v2

@@ III.1 $A(q^{2})$ form factor
Using the numerical solutions of (14) and (26), we numerically solve the integral in (29) as a function of $q^{2}$. In order to compare it with the lattice QCD results, we consider a normalization of our numerical $A(q^{2})$ by $A(0)$ of the lattice data, since this initial value is not fixed in holography. Using an analytical soft-wall model, the authors in Mamo and Zahed (2020) proceeded by choosing $A(0)$ from the lattice data of Shanahan and Detmold (2019a) to fix their corresponding zero momentum squared value and adjust their analytical result to these data. This can be achieved by considering a target value, say $A(0)_{\text{targ}}$ from the lattice, and then a normalized curve given by
@@
such that $A(0)_{\text{norm}}$ coincides with the target value from the lattice.
Here we choose to compare our holographic model with the more up to date Lattice data of Hackett et al. (2024) related to the gluon contribution to $A(q^{2})$. In this work, however, we extend the normalization procedure described above by choosing a range of target $A(0)_{\text{targ}}$ values that comprises the zero momentum values of Shanahan and Detmold (2019a) and Hackett et al. (2024), as well as their error bars. We choose this range as $A(0)_{\text{targ}}=0.4-0.7$. Then, we perform a $\chi^{2}$ minimization procedure as follows. The following $\chi^{2}$ functional form is given by
$\chi^{2}=\sum_{i=2}^{N}\frac{[A(q^{2}_{i})_{\text{norm}}-A(q^{2}_{i})_{\text{lat}})]^{2}}{\sigma^{2}},$ (31)
where $i=2$ corresponds to the second value of both the normalized holographic model and the lattice value, since they are built to be equal at $i=1\,(A(q^{2}=0))$ and hence give $\chi^{2}=0$. $N$ represents the total number of momenta transferred. $\sigma=0.5\,(A_{\text{upper}}-A_{\text{lower}})$ corresponds to the symmetric error associated with the upper and lower points of the error bars of the data from Hackett et al. (2024). Keeping all other parameters in the model fixed, we look for the value of $A(0)_{\text{targ}}$ in $A(q^{2}_{i})_{\text{norm}}$ that minimizes (31). In this context, we find $A(0)_{\text{targ}}=0.628$ to be the best target value found. We point out that this target value was entered as the only degree of freedom used in this minimization procedure. This gives $\chi^{2}_{\text{min}}=32.472$ and $\chi^{2}/\text{ndf}=0.984$. These results give the best-fit curve of our normalized holographic model, as illustrated in Fig. 4.
Figura 4: $A(q^{2})_{\text{norm}}$ form factor adjusted to the gluon contributions of lattice data of Shanahan and Detmold (2019a) and Hackett et al. (2024). The green dots correspond to the gluonic contribution in lattice QCD of Hackett et al. (2024), and the blue curve represents our best-fit normalized holographic form factor AA .The purple curve represents the soft-wall result of Mamo & Zahed Mamo and Zahed (2020), compared to the lattice data of Shanahan and Detmold (2019a) (red dots).
The gluonic radius related to the momentum distribution is given by Polyakov and Schweitzer (2018)
@@ III.2 $D(q^{2})$ form factor
From the numerical integration of (33), one can calculate the form factor $D$ defined by $D(q^{2})\equiv 4\,C(q^{2})$. We perform the same normalization and $\chi^{2}$ minimization procedure for $D(q^{2})$. Similarly to $A(0)_{\text{targ}}$, we choose to restrict the target $D(0)$ in a normalized $D(q^{2})_{\text{norm}}$ to the range $D(0)_{\text{targ}}=(-1.5)-0.0$, in which most of the lattice data from Hackett et al. (2024) lie. The resulting best target value $D(0)_{\text{targ}}$ is found to be $D(0)_{\text{targ}}=-1.192$. The resulting fit to the lattice data of Hackett et al. (2024) is illustrated in Fig. 5. We can notice that we get close to the QCD data of the lattice, but the fit is not as good as that achieved for the form factor $A$, especially for the high momenta transferred $q^{2}$. Indeed, the minimization procedure described above, when applied here to $D(q^{2})_{\text{norm}}$ with a target $D(0)_{\text{targ}}$, gives $\chi^{2}_{\text{min}}=64.219$ with $\chi^{2}/\text{ndf}=4.281$, which quantitatively states that the model deviates a considerable amount from the data considered. In fact, this seems to be a common feature shared between bottom-up holographic models such as the one used in this work and the soft-wall model considered in Mamo and Zahed (2020). However, even though there is this discrepancy between our holographic model and the data considered for comparison, we can still explore the fact that its trend qualitatively follows the gluon lattice data of Shanahan and Detmold (2019a) and Hackett et al. (2024) to study the mechanical properties of the proton, such as the pressure and shear distributions in this model, which we are going to do next.
Figura 5: Holographic $D(q^{2})_{\text{norm}}$ form factor from deformed AdS model (blue curve) compared to gluon lattice data of Hackett et al. (2024) (green dots). The purple curve corresponds to the soft-wall model of Mamo and Zahed (2020), which compares to the lattice data of Shanahan and Detmold (2019a) (red dots).
@@ IV Pressure, Shear, and Energy Distributions
This is because these distributions can be formulated in terms of $\tilde{D}(r)$ according to Polyakov and Schweitzer (2018)
@@
with $m=M_{p}$ the proton mass. For numerical convenience, instead of using the numerical result for $D(q^{2})_{\text{norm}}$, we use a dipole approximation of it, which is well behaved when computing its Fourier transform. We consider
$D(q^{2})=\frac{D_{0}}{\left(1+\frac{q^{2}}{\mathcal{E}^{2}}\right)^{2}},$ (37)
where $D_{0}$ and $\mathcal{E}$ are free parameters. A very good fit to our numerical result $D(q^{2})_{\text{norm}}$ is found with $D_{0}=-1.183$ and $\mathcal{E}=1.184\,\text{GeV}$. The pressure and shear distributions are shown by Figures 7 and 7.
Figura 6: Pressure distribution from the dipole approximation (37), with a repulsive (red) and confining (green) pressure regions.
Figura 7: Shear distribution from the dipole approximation (37).
One can also obtain the mechanical radius associated to those distributions, given by Polyakov and Schweitzer (2018)
$\langle\,r_{\text{mech}}^{2}\rangle=\frac{6}{\mathcal{E}^{2}}=(0.41\,\text{fm})^{2}.$ (38)
These results are comparable with experimental results in Burkert et al. (2018) and lattice QCD ones, as in Shanahan and Detmold (2019b).
One can notice from these predictions that the strong force behaves as a repulsive force in the inner regions of the proton, whereas it behaves as a confining force in its outer regions. These positive and negative pressure regions present in Fig. 7 integrate to $1.93\times 10^{-6}\,\text{GeV}$, which, within the precision of our calculation, is numerically consistent with the von Laue stability condition Laue (1911)
@@
$\epsilon(r)=m\,\left[A(q^{2})-\frac{q^{2}(D(q^{2})+A(q^{2})-2\,B(q^{2}))}{4\,m^{2}}\right]_{\text{FT}},$ (40)
where FT stands for Fourier transform. Here, the $A$ and $D$ form factors represent our normalized ones obtained in Sect. III and, in our case, $B=0$. We also need to obtain an approximation to our numerical $A(q^{2})_{\text{norm}}$ to compute the Fourier transform above, and again consider a dipole parameterization of the form
@@
with $A_{0}$ and $\lambda$ free parameters. The best approximation to our numerical $A(q^{2})_{\text{norm}}$ is found with $A_{0}=0.622$ and $\lambda=1.201\,\text{GeV}$. Hence, the resulting energy density can be computed and is displayed in Fig. 8.
Figura 8: Energy distribution from the dipole approximations (37) and (41).
@@ V Normal and Tangential Forces
Following Polyakov and Schweitzer (2018), we consider a slice of the proton defined in spherical coordinates by $\theta=\pi/2$. The forces acting on the infinitesimal area element $d\textbf{S}=dS_{n}\,\textbf{e}_{n}+dS_{\theta}\,\textbf{e}_{\theta}+dS_{\phi}\,\textbf{e}_{\phi}+$ of the slice are
@@
Figura 9: Normal and tangential forces on a spherical shell of radius $r$ in the nucleon.
@@
Figura 10: Force field visualization of the tangential forces $F_{\theta}$ and $F_{\phi}$ as functions of $r$. The left panel displays the full profile of these forces. The right panel zooms in into the inner region where the sign of the forces changes.
Figura 11: Force field visualization of the normal force $F_{n}(r)$ inside the proton.
@@ VI Conclusions
In the present work, we employ the deformed AdS background holographic model to describe gravitational form factors of the proton and explore some of its mechanical properties based on the $D$ term, focusing on the contribution of the gluon to them. We find a very good agreement between our prediction for the gluonic form factor $A$ and the recent lattice QCD computation performed in Hackett et al. (2024), using a normalization procedure to fit $A(0)$ to the lattice result with a target $A(0)_{\text{targ}}$ that minimizes the $\chi^{2}$ measure of the difference between our model prediction and the lattice data. Its trend is also compared to the previous holographic result of Mamo and Zahed (2020) as well as to the previous lattice data of Shanahan and Detmold (2019a). The same approach is carried out for the gluonic term $D$, resulting, in this case, in a qualitative agreement between our model and the lattice data of Hackett et al. (2024), whose behavior draws a particular shape not fully captured by the holographic model used in this work. However, we remark that those results were obtained numerically in all the steps present in their calculations, in contrast to previous holographic results using the soft-wall model in Abidin and Carlson (2009); Mamo and Zahed (2020), which use analytical solutions to form factors. The general agreement between the numerical results in our work, which are also in the ballpark of the results present in Shanahan and Detmold (2019a); Mamo and Zahed (2020), and the lattice results of Hackett et al. (2024), shows that this model is capable of capturing features present in both the lattice QCD literature and previous holographic models.
@@
From the $D$ term, since its trend follows the behavior present in the lattice results, we can explore how the model describes some mechanical properties of the nucleon. Using a dipole approximation to our numerical result, which is better suited for the calculation of the Fourier transform of $D(r)$ and $A(r)$, we can compute the pressure and shear distributions inside the proton, which balance each other to keep the system stable. This is supported by the small value of the integral $\int r^{2}\,p(r)\,dr=1.93\times 10^{-6}\,\text{GeV}$ in our approximation, which should be zero in the exact case. We are also able to evaluate the energy density in the proton using a dipole approximation to $A(q^{2})_{\text{norm}}$. The results follow similar trends present in the literature, e.g., such as those in Lorcé et al. (2019); Polyakov and Schweitzer (2018); Shanahan and Detmold (2019b); Burkert et al. (2018, 2021); Mamo and Zahed (2020, 2021). However, the mechanical radius associated with these distributions is smaller than the ones found in these works. From pressure and shear distributions, internal forces are also investigated, presenting a similar behavior appearing in this literature.
We remark that even though the calculations performed in this work follow a similar receipt in holographic modeling as the ones performed in Abidin and Carlson (2009); Mamo and Zahed (2020) and have comparable results, they are obtained in a fundamentally different manner on theoretical grounds. The approach used here within the deformed AdS model considers a direct deformation of the anti-de Sitter space by a dilaton-like term (2) present in its metric (1), as opposed to the soft-wall model we briefly discussed in Appendix A. This simple change makes a real difference, especially in the description of fermions. Beyond fermions, the graviton solutions in the soft-wall model are analytical, while in the deformed AdS model they are numerical, which consequently leads to numerical form factors as the ones shown in Sect. III. This casts the whole approach using the deformed AdS model fundamentally different from the one using the soft-wall model, resulting in a new prediction of the gravitational form factors parameterizing the proton matrix elements of the QCD energy-momentum tensor, as well as the resulting mechanical properties associated with it.